OpenStudy (anonymous):

how do i solve x+1 2x-2 ___ = ____ 5 6

OpenStudy (anonymous):

I just need to know where to start from and I can do the rest from there.. will stillgive medals though

OpenStudy (cwrw238):

multiply both sides of the equation by the LCM of 5 and 6 which is 30

OpenStudy (arbershabani97):

you just multiply x+1 with 6 and 2x-2 with 5 6x+6=10x-10 -4x=-4 x=1

OpenStudy (anonymous):

No x is equal to 4 just by looking at it..but im not sure how to mathematically get it

OpenStudy (anonymous):

You would have to multiply everything on both sides by 30 (Least common factor of 5 and 6) and then isolate x.

OpenStudy (anonymous):

Like cwrw238 said

OpenStudy (anonymous):

ohh okay

OpenStudy (cwrw238):

6x + 6 = 10x - 10 16 = 4x

OpenStudy (arbershabani97):

you just multiply x+1 with 6 and 2x-2 with 5 6x+6=10x-10 -4x=-16 x=4

OpenStudy (cwrw238):